Please help me solve: (-3,__) y=4-3x
Answer by nerdybill(7384)

(-3,__) y=4-3x
.
(-3,___) gives you the 'x-value" of an (x,y) coordinate pair.
.
So, now that you know x=-3, plug into the equation to find y:
y=4-3x
y=4-3(-3)
y=4-(-9)
y=4+9
y=13
.
Answer: (-3, 13)